OpenStudy (anonymous):

dose anyone really get how to find the slopes of straight lines? i just don't get it!

OpenStudy (anonymous):

i believe you use y=mx+b

OpenStudy (anonymous):

pick 2 known points. use M=(y2-y1)/((x2-x1)

OpenStudy (anonymous):

what is the equation

OpenStudy (anonymous):

then sub in the values of your known point into the fourmula y-y=m(x-x)

OpenStudy (anonymous):

ok i will ask my parents for some more help....but for now i g2g...thanks for the help!!!

OpenStudy (anonymous):

lol this should hellp http://tutorial.math.lamar.edu/Classes/Alg/Graphing_Functions.aspx

OpenStudy (amistre64):

you find the slope between two points by measuring how far you have to rise and how far you have to run to get there. like stairs

OpenStudy (amistre64):

slope is defined in math as a ration between: the change in the distance of you "y" values ------------------------------------- the change in the distance of you "x" values

OpenStudy (amistre64):

change in distance and direction :)

OpenStudy (amistre64):

some slope go up... some slope go down..... some slope is flat..... some slope is to far to see the top ;)